OpenStudy (anonymous):

If the sides of a square are lengthened by 5m, the area becomes 196cm^2. Find the length of a side of the original square.

OpenStudy (anonymous):

\[A=s^{2}\] for a regular square For this square, \[A=(s+5)^{2}\] so \[196=(s+5)^{2}\]

OpenStudy (mertsj):

(x+5)^2=196 x+5=14 x=9

OpenStudy (anonymous):

So if s or x = 9, then to find the length of a side of the original square I need to plug 9 into A=s^2, which would be A=9^2, A=81?

OpenStudy (mertsj):

No. 9 is the side of the original square. Notice the equation in relation to the problem. It says the side of the original square was increased by 5 so x+5 means the original side with 5 added. So x is the original side.

OpenStudy (anonymous):

Ok, I understand. Thank you.